OpenStudy (anonymous):

150 people live in a village. Of these, 60 are Adults, 90 are children. Calculate, in it's lowest terms, the ratio of adults to children.

OpenStudy (anonymous):

the ratio of adults to children is : total adults / total children = 60/90 = 2/3